# Critical behaviour of the fully packed loop-$O(n)$ model on planar triangulations

### v2

@@ Abstract
We study the fully packed loop-$O(n)$ model on planar triangulations. This model is also bijectively equivalent to the Fortuin–Kasteleyn model of planar maps with parameter $q\in(0,4)$ at its self-dual point. These have been traditionally studied using either techniques from analytic combinatorics (based in particular on the gasket decomposition of Borot, Bouttier and Guitter [BBG12a]) or probabilistic arguments (based on Sheffield’s “hamburger-cheeseburger” bijection [sheffield2016quantum]). In this paper we establish a dictionary relating quantities of interest in both approaches. This has several consequences. First, we derive an exact expression for the partition function of the fully packed loop-$O(n)$ model on triangulations, as a function of the outer boundary length. This confirms predictions by Gaudin and Kostov [gaudin1989n]. In particular, this model exhibits critical behaviour, in the sense that the partition function exhibits a power-law decay characteristic of the critical regime at this self-dual point. Finally, we derive precise asymptotics for geometric features of the FK model of planar maps when $0<q<4$, such as the exact tail behaviour of the perimeters of clusters and loops. This sharpens previous results of [berestycki2017critical] and [gwynne2019scaling]. A key step is to use the above dictionary and the probabilistic results to justify rigorously an ansatz commonly assumed in the analytic combinatorics literature.
@@ 1 Introduction
Planar maps (i.e., proper embeddings of graphs in the $2$-sphere, considered up to orientation-preserving homeomorphisms) are a central topic not only in combinatorics but also in probability and mathematical physics due to their conjectured links with Liouville quantum gravity (LQG). In that context, planar maps can be thought of as canonical discretisations of the “random surfaces” which are at the core of Polyakov’s original formulation of LQG [polyakov1981quantum]. Indeed, in order to describe the gravitational action when gravity is coupled to a matter field, the so-called DDK Ansatz ([David, DistlerKawai]) implies that this can be equivalently described by considering the scaling limit of random planar maps decorated by models of statistical mechanics at their critical point.
In this paper we will consider two such models, which turn out to be bijectively related. We start with the Fortuin–Kasteleyn percolation model with parameter $q\in(0,4)$. In this model, we sample a pair $(M,\Omega)$ where $M$ is a (rooted) planar map with a fixed number of edges (say $k$ edges), and $\Omega$ is a subset of edges of $M$. In general, the probability of sampling a given decorated map $(\mathfrak{m},\omega)$ depends on an extra percolation parameter $p_{0}\in(0,1)$ and a volume weight $s$ and is proportional to
$\mathbb{P}((M,\Omega)=(\mathfrak{m},\omega))\,\propto\,\left(\frac{p_{0}}{1-p_{0}}\right)^{\mathsf{o}(\omega)}q^{\mathsf{cc}(\omega)}s^{\mathsf{v}(\mathfrak{m})},$ (1.1)
where $\mathsf{o}(\omega)$ and $\mathsf{cc}(\omega)$ are the number of edges and connected components of $\omega$, and $\mathsf{v}(\mathfrak{m})$ is the number of vertices of $\mathfrak{m}$. In particular, conditional on the planar map $M=\mathfrak{m}$, the edge configuration $\Omega$ is sampled as an FK(q)(q)-percolation configuration on $\mathfrak{m}$ (see, e.g., [duminil2013parafermionic] for a general introduction to this model). On such a decorated planar map $(\mathfrak{m},\omega)$ there is a natural duality operation $(\mathfrak{m},\omega)\to(\mathfrak{m}^{\dagger},\omega^{\dagger})$. Requiring that the law (1.1) is invariant under this duality imposes that $p_{0}=\sqrt{q}/(1+\sqrt{q})$ and $s=1/\sqrt{q}$, as can be checked using Euler’s formula. In that case the law (1.1) reduces to
@@
where $\mathsf{loops}(\mathfrak{m},\omega)$ is the set of loops separating $\omega$ and its dual in $\mathfrak{m}$. This is called the self-dual FK(q)(q)-weighted planar map model. In the self-dual case, one can also let the number of edges $k\to\infty$ to obtain, as a local limit of the above, an infinite FK(q)(q)-weighted planar map model [sheffield2016quantum, chen2017basic].
@@
On the one hand, from the perspective of the self-dual FK model (1.2), criticality (if it indeed holds) would agree with the fact that, in common with many planar models of statistical mechanics, the self-dual point coincides with its transition point (see for instance [grimmett2018probability] for some background discussion). In particular, such a statement is the content of a celebrated result of Beffara and Duminil-Copin [beffara2012self] for the random cluster model (equivalently the Fortuin–Kasteleyn percolation model) on the square lattice and for $q\geq 1$.
@@
On the hexagonal lattice, the loop-$O(n)$ model (1.5) is predicted to have a rich phase diagram (see e.g. [peled2019lectures, Section 3]). Notably, it was predicted by Nienhuis [Nienhuis82, Nienhuis84] that there is a critical line, given by
@@
for $n\in(0,2]$ (and in fact $n\in[-2,2]$ in the physics literature) separating the sub-critical phase ($z<z_{c}$) where the loops exhibit exponential decay and the critical phase ($z\geq z_{c}$) where the decay should only be polynomial. In fact, for fixed $n$, there should be two “critical” regimes called dilute and dense respectively, according as $z=z_{c}(n)$ or $z>z_{c}(n)$, where the model is expected to have different (but both conformally-invariant) scaling limits. See [kager2004guide, Section 5.6] for a statement of this conjecture. Although this remains a famous open problem in the field, we also stress that spectacular progress has been made in various regions of the phase diagram – we refer to the recent paper [glazman2025planar] for the state of the art and an important breakthrough in this direction. It is not altogether clear whether the fully-packed case on the hexagonal lattice should correspond to a third phase with a distinct conformally invariant scaling limit or should be in the same universality class as the dense phase (see [blote1994fully] for some discussion). Nonrigorous renormalisation arguments have suggested that the fully-packed case corresponds to an unstable fixed (critical) point, although on the square lattice, according to some predictions [blote2012completely] the fully packed model appears to behave as a dense phase. In the setting of planar maps, there seems to be a consensus that the fully-packed case belongs to the same universality class as the dense case and our results below support this. It would be interesting to study the fully packed model further on the hexagonal and square lattices, since typically, one expects the same behaviour on a fixed lattice and a random one when there is universality and/or conformal invariance.
@@
On the other hand, a classical approach to the loop-$O(n)$ model is to make use of the spatial Markov property of the model. This results in the so-called gasket decomposition which was in particular used by Borot, Bouttier and Guitter [BBG12, BBG12a, BBG12b] to establish the phase diagram of the loop-$O(n)$ model. Building on powerful tools of analytic combinatorics they derived and solved (assuming a certain natural ansatz) an equation for the resolvent of the partition function of the model, which yields fine asymptotics. Such an approach has been made fully rigorous (including the proof of the above ansatz) by works of Budd and Chen [budd2019peeling], in the case of the so-called rigid model on quadrangulations (where loops are constrained to enter and leave a given quadrangle through opposite edges). From this, Chen, Curien and Maillard [ChenCurienMaillard] deduced some scaling limit results for the perimeter cascade of $O(n)$ loops, and Aïdékon, Da Silva and Hu [aidekon2024scaling] proved the scaling limit for the volume of such rigid loop-$O(n)$ quadrangulations. A similar approach was also used by Borot, Duplantier and Guitter [BBD16] to determine the nesting statistics in the bending energy variant of the $O(n)$ model, where loops can bend inside a quadrangle at a given energetic cost.
However, we emphasise that the aforementioned ansatz (and thus, the phase diagram of the model) has so far only been rigorously established in the rigid case [budd2019peeling]. The proof relies on special symmetries of bipartite maps and cannot be directly extended to the case of triangulations. We comment on the exact nature of the missing step in Section 2.3.2.
@@
Observe that the form of the asymptotic (1.7) is characteristic of the critical (in fact, more precisely, non-generic critical according to the terminology of [BBG12]) phase of the loop-$O(n)$ model, with the exponent $2-\theta$ interpolating between $3/2$ and $2$ (see e.g. [BBG12, Section 5]). More precisely, it lies in the so-called dense phase of the model, where loops are conjectured to be non-simple and to touch each other in the scaling limit.
Our second main result determines the exact tail exponents for the size of typical loops and clusters in the original FK(q)(q)-weighted planar map model. This sharpens the main theorem of [berestycki2017critical] by identifying the previously implicit $\ell^{o(1)}$ factor as a true constant (a similar result can also be deduced from [gwynne2019scaling], with the $\ell^{o(1)}$ term identified as a slowly varying function). We denote by $\mathfrak{L}$ and $\mathfrak{K}$ the typical loop and (filled-in) cluster in the infinite FK(q)(q) planar map. These correspond to the local limit of a uniformly chosen loop or (filled-in) cluster in a finite FK(q)(q) map of size $k$, and then sending $k$ to infinity, see [berestycki2017critical, Theorem 1.1]. We define the perimeter $|\mathfrak{L}|$ of the loop $\mathfrak{L}$ as the number of triangles it crosses, and the perimeter $|\partial\mathfrak{K}|$ of the cluster as the degree of its external face, see Section 3.1 for more precise definitions.
@@
The paper is organised as follows. Section 2 is dedicated to preliminaries on self-dual FK planar maps, the $O(n)$ model and the Mullin–Bernardi–Sheffield bijection. In Section 3, we make some connections between Sheffield’s hamburger-cheeseburger walk and the gasket decomposition of Borot, Bouttier and Guitter. In particular, this allows us to prove an ansatz on the cut of the associated resolvent function using solely hamburger-cheeseburger arguments: we will derive an explicit formula for the right endpoint $\gamma_{+}$ of the cut. We will then use this information in Section 4, where we solve the resolvent equation for the fully packed $O(n)$ model on triangulations, thus proving 1.1. Finally, in Section 5, we come back to the hamburger-cheeseburger model and deduce 1.3 by plugging back the information on the partition function.
@@
We stress that this dictionary has been used in the concurrent work [da2025scaling], which concerns the critical case $q=4$ (equivalently, $n=2$). In that case, the authors prove the analogue of 1.1 by directing solving the gasket decomposition equation (without any extra input) and then deduce the scaling limit of FK(4)(4)-weighted planar maps in the peanosphere sense, resolving the remaining open regime in Sheffield’s work [sheffield2016quantum].

### v3

# The self-dual point of Fortuin–Kasteleyn planar maps is critical

@@ Abstract
We study the Fortuin–Kasteleyn model of planar maps with parameter $q\in(0,4)$ at and away from its self-dual point. This model is also bijectively equivalent to the fully packed (bicoloured) loop-$O(n)$ model on planar triangulations. These have been traditionally studied using either techniques from analytic combinatorics (based in particular on the gasket decomposition of Borot, Bouttier and Guitter [BBG12a]) or probabilistic arguments (based on Sheffield’s “hamburger-cheeseburger” bijection [SHE16]). In this paper we establish a dictionary relating quantities of interest in both approaches. This has several consequences. First, we derive an exact expression for the partition function of the fully packed (colour-symmetric) loop-$O(n)$ model on triangulations, as a function of the outer boundary length. This confirms predictions by Gaudin and Kostov [GK89]. In particular, this model exhibits critical behaviour, in the sense that the partition function exhibits a power-law decay characteristic of the critical regime at this self-dual point. Secondly, we derive precise asymptotics for geometric features of the self-dual Fortuin–Kasteleyn model of planar maps when $0<q<4$, such as the exact polynomial tail behaviour of the perimeters of clusters and loops. This sharpens previous results of [BLR17] and [GMS19]. Finally, we prove that Fortuin–Kasteleyn maps undergo a sharp phase transition at the self-dual point in the sense that, away from the self-dual point, cluster sizes decay exponentially. This mirrors the celebrated result by Beffara and Duminil-Copin for $q\geq 1$ on a fixed lattice [BD12] and constitutes the first result establishing that the self-dual point of Fortuin–Kasteleyn maps is the critical point. Along the way and as a key step of our proof, we use the above dictionary and the probabilistic results to justify rigorously an ansatz commonly assumed in the analytic combinatorics literature.
@@ 1 Introduction
Planar maps (i.e., proper embeddings of graphs in the $2$-sphere, considered up to orientation-preserving homeomorphisms) are a central topic not only in combinatorics but also in probability and mathematical physics due to their conjectured links with Liouville quantum gravity (LQG). In that context, planar maps can be thought of as canonical discretisations of the “random surfaces” which are at the core of Polyakov’s original formulation of LQG [POL81]. Indeed, in order to describe the gravitational action when gravity is coupled to a matter field, the so-called DDK Ansatz ([DAV88, DK89]) implies that this can be equivalently described by considering the scaling limit of random planar maps decorated by models of statistical mechanics at their critical point.
In this paper we will consider two such models, which turn out to be bijectively related. We start with the Fortuin–Kasteleyn percolation model with parameter $q\in(0,4)$. In this model, we sample a pair $(M,\Omega)$ where $M$ is a (rooted) planar map, and $\Omega$ is a subset of edges of $M$. In general, the probability of sampling a given decorated map $(\mathfrak{m},\omega)$ depends on an extra percolation parameter $p_{0}\in(0,1)$ and a volume weight $s$ and is proportional to
$\mathbb{P}((M,\Omega)=(\mathfrak{m},\omega))\,\propto\,\left(\frac{p_{0}}{1-p_{0}}\right)^{\mathsf{o}(\omega)}q^{\mathsf{cc}(\omega)}(1/\sqrt{q})^{\mathsf{v}(\mathfrak{m})}s^{2\mathsf{e}(\mathfrak{m})},$ (1.1)
where $\mathsf{o}(\omega)$ and $\mathsf{cc}(\omega)$ are the number of edges and connected components of $\omega$, and $\mathsf{v}(\mathfrak{m})$ and $\mathsf{e}(\mathfrak{m})$ are the number of vertices and edges of $\mathfrak{m}$. The term $(1/\sqrt{q})^{\mathsf{v}(\mathfrak{m})}$ is for later convenience (see [BBG12b, Equation (2.10)]). We will mainly be focusing on a the model with fixed number of edges (say $k$ edges), which then becomes
$\mathbb{P}((M,\Omega)=(\mathfrak{m},\omega))\,\propto\,\left(\frac{p_{0}}{1-p_{0}}\right)^{\mathsf{o}(\omega)}q^{\mathsf{cc}(\omega)}(1/\sqrt{q})^{\mathsf{v}(\mathfrak{m})}.$ (1.2)
In particular, conditional on the planar map $M=\mathfrak{m}$, the edge configuration $\Omega$ is sampled as an FK(q)(q)-percolation configuration on $\mathfrak{m}$ (see, e.g., [DUM13] for a general introduction to this model). On such a decorated planar map $(\mathfrak{m},\omega)$ there is a natural duality operation $(\mathfrak{m},\omega)\to(\mathfrak{m}^{\dagger},\omega^{\dagger})$. Requiring that the law (1.2) is invariant under this duality imposes that $p_{0}=p_{\mathrm{sd}}(q):=\sqrt{q}/(1+\sqrt{q})$, as can be checked using Euler’s formula. In that case the law (1.2) reduces to
@@
where $\mathsf{loops}(\mathfrak{m},\omega)$ is the set of loops separating $\omega$ and its dual in $\mathfrak{m}$. This is called the self-dual FK(q)(q)-weighted planar map model. In the self-dual case, one can also let the number of edges $k\to\infty$ to obtain, as a local limit of the above, an infinite FK(q)(q)-weighted planar map model [SHE16, CHE17].
@@
On the one hand, from the perspective of the self-dual FK model (1.3), criticality (if it indeed holds) would agree with the fact that, in common with many planar models of statistical mechanics, the self-dual point coincides with its transition point (see for instance [GRI18] for some background discussion). In particular, such a statement is the content of a celebrated result of Beffara and Duminil-Copin [BD12] for the random cluster model (equivalently the Fortuin–Kasteleyn percolation model) on the square lattice and for $q\geq 1$.
@@
On the hexagonal lattice, the loop-$O(n)$ model (1.6) is predicted to have a rich phase diagram (see e.g. [PS19, Section 3]). Notably, it was predicted by Nienhuis [NIE82, NIE84] that there is a critical line, given by
@@
for $n\in(0,2]$ (and in fact $n\in[-2,2]$ in the physics literature) separating the sub-critical phase ($z<z_{c}$) where the loops exhibit exponential decay and the critical phase ($z\geq z_{c}$) where the decay should only be polynomial. In fact, for fixed $n$, there should be two “critical” regimes called dilute and dense respectively, according as $z=z_{c}(n)$ or $z>z_{c}(n)$, where the model is expected to have different (but both conformally-invariant) scaling limits. See [KN04, Section 5.6] for a statement of this conjecture. Although this remains a famous open problem in the field, we also stress that spectacular progress has been made in various regions of the phase diagram – we refer to the recent paper [GHZ25] for the state of the art and an important breakthrough in this direction. It is not altogether clear whether the fully-packed case on the hexagonal lattice should correspond to a third phase with a distinct conformally invariant scaling limit or should be in the same universality class as the dense phase (see [BN94] for some discussion). Nonrigorous renormalisation arguments have suggested that the fully-packed case corresponds to an unstable fixed (critical) point, although on the square lattice, according to some predictions [BWG12] the fully packed model appears to behave as a dense phase. In the setting of planar maps, there seems to be a consensus that the fully-packed case belongs to the same universality class as the dense case and our results below support this. It would be interesting to study the fully packed model further on the hexagonal and square lattices, since typically, one expects the same behaviour on a fixed lattice and a random one when there is universality and/or conformal invariance.
@@
On the other hand, a classical approach to the loop-$O(n)$ model is to make use of the spatial Markov property of the model. This results in the so-called gasket decomposition which was in particular used by Borot, Bouttier and Guitter [BBG12b, BBG12a, BBG12c] to establish the phase diagram of the loop-$O(n)$ model. Building on powerful tools of analytic combinatorics they derived and solved (assuming a certain natural ansatz) an equation for the resolvent of the partition function of the model, which yields fine asymptotics. Such an approach has been made fully rigorous (including the proof of the above ansatz) by works of Budd and Chen [BC19], in the case of the so-called rigid model on quadrangulations (where loops are constrained to enter and leave a given quadrangle through opposite edges). From this, Chen, Curien and Maillard [CCM20] deduced some scaling limit results for the perimeter cascade of $O(n)$ loops, and Aïdékon, Da Silva and Hu [ADH26] proved the scaling limit for the volume of such rigid loop-$O(n)$ quadrangulations. A similar approach was also used by Borot, Duplantier and Guitter [BBD23] to determine the nesting statistics in the bending energy variant of the $O(n)$ model, where loops can bend inside a quadrangle at a given energetic cost.
However, we emphasise that the aforementioned ansatz (and thus, the phase diagram of the model) has so far only been rigorously established in the rigid case [BC19]. The proof relies on special symmetries of bipartite maps and cannot be directly extended to the case of triangulations. We comment on the exact nature of the missing step in Section 2.3.2.
@@
Observe that the form of the asymptotic (1.8) is characteristic of the critical (in fact, more precisely, non-generic critical according to the terminology of [BBG12b]) phase of the loop-$O(n)$ model, with the exponent $2-\theta$ interpolating between $3/2$ and $2$ (see e.g. [BBG12b, Section 5]). More precisely, it lies in the so-called dense phase of the model, where loops are conjectured to be non-simple and to touch each other in the scaling limit.
Our second main result determines the exact tail exponents for the size of typical loops and clusters in the original FK(q)(q)-weighted planar map model. This sharpens the main theorem of [BLR17] by identifying the previously implicit $\ell^{o(1)}$ factor as a true constant (a similar result can also be deduced from [GMS19], with the $\ell^{o(1)}$ term identified as a slowly varying function). We denote by $\mathfrak{L}$ and $\mathfrak{K}$ the typical loop and (filled-in) cluster in the infinite FK(q)(q) planar map. These correspond to the local limit of a uniformly chosen loop or (filled-in) cluster in a finite FK(q)(q) map of size $k$, and then sending $k$ to infinity, see [BLR17, Theorem 1.1]. We define the perimeter $|\mathfrak{L}|$ of the loop $\mathfrak{L}$ as the number of triangles it crosses, and the perimeter $|\partial\mathfrak{K}|$ of the cluster as the degree of its external face, see Section 3.1 for more precise definitions.
@@
The last main result of this paper concerns the Fortuin–Kasteleyn model away from self-duality. For this result we work under the free model (1.1) with arbitrary number of edges. Recall the self-dual parameter $p_{\mathrm{sd}}(q)=\sqrt{q}/(1+\sqrt{q})$, and define the two regions
$\mathcal{R}_{q}:=\Big\{(s,p_{0})\in\mathbb{R}_{+}^{*}\;\Big|\;s\leq x_{c},\,s\sqrt{\frac{p_{0}}{\sqrt{q}(1-p_{0})}}\leq x_{c}\Big\}\quad\text{and}\quad\mathcal{S}_{q}:=\mathcal{R}_{q}\setminus\{(x_{c},p_{\mathrm{sd}}(q))\}.$ (1.9)
Although it is not clear at this point, we will see in Section 2.3.1 that the partition function of the FK($q$) model (1.1) with parameters $(s,p_{0})\in\mathcal{R}_{q}$ is finite. Since the partition function of this model is finite, we may therefore normalise the weights (1.1) to get a probability distribution $\mathbb{P}_{(s,p_{0})}$; we then call $\mathfrak{K}$ the corresponding (filled-in) random cluster of the root edge, which may be of any colour $i=1,2$, corresponding to being primal or dual. (Note that in contrast with the notation used in 1.3, here the law $\mathbb{P}_{(s,p_{0})}$ under which this random variable is considered is not that of any local limit.) In addition, we interpret $\mathcal{S}_{q}$ as a subcritical region of the FK($q$) model (1.1), since the condition defining $\mathcal{S}_{q}$ in (1.9) guarantees that the model stays away from its self-dual point $(x_{c},p_{\mathrm{sd}}(q))$. We note that since $\frac{p_{\mathrm{sd}}(q)}{1-p_{\mathrm{sd}}(q)}=\sqrt{q}$, the region $\mathcal{R}_{q}$ contains in particular all pairs $(s,p_{0})$ with $s\leq x_{c}$ and $p_{0}\leq p_{\mathrm{sd}}(q)$ (in addition, note that the region $\mathcal{S}_{q}$ also contains values of $(s,p_{0})$ with $p_{0}>p_{\mathrm{sd}}(q)$ by taking $s$ small enough). We let $K_{\ell}^{(i)}$, $i=1,2$, be the partition function of Fortuin–Kasteleyn maps such that the root cluster has colour $i$ and perimeter $\ell$. Our final main theorem shows the exponential decay of $K_{\ell}^{(i)}$ for $i=1,2$ in the subcritical phase $\mathcal{S}_{q}$.
Theorem 1.4 (Exponential decay of $K_{\ell}^{(i)}$).
Consider the $\mathrm{FK}(q)$ model described in (1.1) and assume that $(s,p_{0})\in\mathcal{S}_{q}$. Then, for all $i=1,2$ there exists a constant $C_{i}>0$ and $\gamma_{i}\in(0,1)$ such that for all $\ell\geq 1$,
$K_{\ell}^{(i)}\leq C_{i}\gamma_{i}^{\ell}.$
In contrast, when $(s,p_{0})\in\mathcal{R}_{q}\setminus\mathcal{S}_{q}$, the partition functions $K_{\ell}^{(i)}$, $i=1,2$, decay polynomially: more specifically, there is a constant $C>0$ such that
$K_{\ell}^{(1)}=K_{\ell}^{(2)}\sim C\ell^{2\theta-4}\quad\text{as }\ell\to\infty,$
with $\theta$ given by (1.7).
As an immediate consequence of this result we deduce the following: if $(s,p_{0})\in\mathcal{S}_{q}$ then $\mathbb{P}_{(s,p_{0})}(|\partial\mathfrak{K}|=\ell)$ decays exponentially fast as $\ell\to\infty$, whereas it decays polynomially fast (with the stated exponent) when $(s,p_{0})\in\mathcal{R}_{q}\setminus\mathcal{S}_{q}$. 1.4 is thus a statement which shows that the Fortuin–Kasteleyn model of planar maps undergoes a phase transition at the self-dual point, and that this phase transition is furthermore sharp in the language of percolation: while, at the self-dual point, cluster sizes decay polynomially, away from self-duality they are exponentially small. On the square lattice, such a result was first established (for $q\geq 1$) by Beffara and Duminil-Copin [BD12] (see also [DRT19]). To the best of our knowledge, our paper provides the first result showing such a phase transition (and sharpness) for the FK($q$) model on planar maps.
There is however an interesting difference with the case of the infinite planar square lattice. In that case, if $p$ is chosen away from the self-dual point, it is well known that there is either percolation of the primal or dual configuration. By contrast, note that in Theorem 1.4 above, if $(s,p_{0})\in\mathcal{S}_{q}$ then there is exponential decay of both $K^{(1)}_{\ell}$ and $K^{(2)}_{\ell}$, in particular irrespective of whether $p_{0}<p_{\mathrm{sd}}(q)$ or $p_{0}>p_{\mathrm{sd}}(q)$. This suggests that there may not exist a canonical way to make sense of a local limit of these maps away from the self-dual point, even when one imposes the colour of $\mathfrak{K}$ to be the “dominant” one (i.e. primal when $p_{0}>p_{\mathrm{sd}}(q)$ and dual when $p_{0}<p_{\mathrm{sd}}(q)$).
Finally, we stress that (1.4) can actually be phrased in terms of the bicoloured loop-$O(n)$ model presented in more detail in Section 2.3. In that framework, the region $\mathcal{S}_{q}$ corresponds to the set of edge parameters that satisfy $x_{1},x_{2}\leq x_{c}$ and $(x_{1},x_{2})\neq(x_{c},x_{c})$. See Section 2.3.1 and the proof in Section 6 for more details.
@@
The paper is organised as follows. Section 2 is dedicated to preliminaries on self-dual FK planar maps, the $O(n)$ model and the Mullin–Bernardi–Sheffield bijection. In Section 3, we make some connections between Sheffield’s hamburger-cheeseburger walk and the gasket decomposition of Borot, Bouttier and Guitter. In particular, this allows us to prove an ansatz on the cut of the associated resolvent function using solely hamburger-cheeseburger arguments: we will derive an explicit formula for the right endpoint $\gamma_{+}$ of the cut. We will then use this information in Section 4, where we solve the resolvent equation for the fully packed $O(n)$ model on triangulations, thus proving 1.1. Then, in Section 5, we come back to the hamburger-cheeseburger model and deduce 1.3 by plugging back the information on the partition function. Finally, we prove 1.4 in Section 6.
@@
We stress that this dictionary has been used in the concurrent work [DHP+25], which concerns the critical case $q=4$ (equivalently, $n=2$). In that case, the authors prove the analogue of 1.1 by directing solving the gasket decomposition equation (without any extra input) and then deduce the scaling limit of FK(4)(4)-weighted planar maps in the peanosphere sense, resolving the remaining open regime in Sheffield’s work [SHE16].